### v1

# Anton D. Utyushev,a Vadim I. Zakomirnyi,b Alexey A. Shcherbakov,a Ilia L. Rasskazovc and Alexander Morozd

[…]
[Uncaptioned image] Intuitive understanding of extinction of small particles in absorbing and active host media within the MLWA
Anton D. Utyushev,a Vadim I. Zakomirnyi,b Alexey A. Shcherbakov,a Ilia L. Rasskazovc and Alexander Morozd
[Uncaptioned image] In an absorbing or an active host medium characterized by a complex refractive index $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$, our previously developed modified dipole long-wave approximation (MLWA) is shown to essentially overly with the exact Mie theory results for spherical nanoparticle with radius $a\lesssim 25$ nm ($a\lesssim 20$ nm) in the case of Ag and Au (Al and Mg) nanoparticles. The agreement for Au and Ag (Al and Mg) nanoparticles, slightly better in the case of Au than Ag, continues to be acceptable up to $a\sim 50$ nm ($a\sim 40$ nm), and can be used, at least qualitatively, up to $a\sim 70$ nm ($a\sim 50$ nm) correspondingly. A first order analytic perturbation theory (PT) in a normalized extinction coefficient, $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$, around a nonabsorbing host is developed within the dipole MLWA and its properties are investigated. It is shown that, in a suitable parameter range, the PT can reliably capture the effect of host absorption or gain on the extinction efficiency of various plasmonic nanoparticles.
footnotetext: a School of Physics and Engineering, ITMO University, 197101, Saint-Petersburg, Russiafootnotetext: b Beckman Institute for Advanced Science and Technology, University of Illinois at Urbana-Champaign, Urbana, IL 61801, USAfootnotetext: c SunDensity Inc., Rochester, NY 14604, USA footnotetext: d Wave-scattering.com footnotetext:  Corresponding author: ad.utyushev@gmail.comfootnotetext: † Electronic Supplementary Information (ESI) available: See DOI:
1 Introduction
Electromagnetic scattering in an absorbing host characterized by a complex refractive index $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$ (where $n_{2}^{\prime}$ and $n_{2}^{\prime\prime}$ are real) has more than fifty years old history. The traditional scattering theory neglects the host dissipation and gain 1, 2, because those cases imply either vanishing or infinite scattering wave at the spatial infinity. Once wave number $k=k^{\prime}+{\rm i}k^{\prime\prime}$ is a complex number, conventional expressions for cross sections cannot be straightforwardly extended for $k^{\prime\prime}\neq 0$, because the expressions yield cross sections as complex quantities. Not surprisingly, the history of scattering in an absorbing host is filled in with a number of controversies 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19. This is probably why in classical textbooks it is only fleetingly mentioned in Section 12.1.3 of ref. 2. Already the definition of an incident intensity is not straightforward, as the field incident on a particle is different at different points on the particle as a result of an absorbing medium 4, 5. In an absorbing or an active unbounded host, the imaginary part $k^{\prime\prime}$ of the complex wave number $k=k^{\prime}+ik^{\prime\prime}=2\pi n_{2}/\lambda_{0}$, where $\lambda_{0}$ is the vacuum wavelength, is nonzero, $k^{\prime\prime}\neq 0$. A great deal of effort was required to arrive at suitable definitions of cross sections for $k^{\prime\prime}\neq 0$.
In contrast to the conventional case of $k^{\prime\prime}=0$, two sets of cross sections are commonly used: inherent and apparent. The former is obtained by performing surface integrals of corresponding Poynting vectors over the particle surface. The approach was developed a long time ago 4, 5, but currently accepted expressions for the inherent cross-sections were not presented until 1999 7, 8, 9, 10. The history of apparent cross sections, which are operationally defined in the far field, began with the optical theorem of Bohren and Gilra 6. Here, too, it took a long time, until 2007, to arrive at definitive expressions for the apparent extinction cross-section cross section, $C_{\rm ext}$, and the so-called apparent scattering cross section, $C_{\rm sca}$ 11, 12, 13, 14. A curious feature of the apparent extinction cross-section in an absorbing host is that it can be negative 20, which is neither an artifact of numerical simulations nor it violates any physical law. The apparent extinction cross-section quantifies the difference in the readings of a forward-scattering detector taken with and without the particle. If the surrounding medium is absorbing, the presence of the particle can in fact make the detector signal stronger, thereby implying a negative extinction cross-section. There is no violation of the energy conservation law, since in this case the extinction cross-section is not used to quantify the energy budget of a finite volume encompassing the particle 20. Another curiosity of apparent cross-sections in an absorbing host is that an intrinsic definition of an apparent absorption cross-sections, $C_{\rm abs}$, is still missing. The difficulty lies in that the very presence of a particle necessarily modifies the near field around the particle. The latter may be the cause of an additional absorption in the host medium outside the particle compared to what is happening in the absence of the particle 8, 9. The critical point is that because the field is disturbed by the particle, additional absorption may be realized in the medium external to the particle 21.
First-principles far-field computations based on the general Lorenz–Mie theory showed that increasing absorption in the host medium broadens and suppresses plasmon resonances in the apparent extinction and effective scattering efficiency factors and suppresses resonance features in the phase function 14. At the same time, using inherent cross-sections it was shown that a surrounding lossy medium can narrow the plasmonic resonances, as well as increase their amplitude dramatically 15, 17. Given this recent revival of interest in the role of the host dissipation or gain on extinction cross sections with conflicting predictions of it either enhancing or reducing the localized surface plasmonic resonance (LSPR) of small particles 14, 15, 17, 18, 19, the focus of present work is on the apparent extinction cross-section $C_{\rm ext}$ in the so-called modified long-wave approximation (MLWA) 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19. The MLWA 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19, which can be viewed as a next-order approximation beyond the Rayleigh limit, is known to overcome a number of severe deficiencies of the Rayleigh limit (see eqn (2) below) and, at least for nonabsorbing hosts, be surprisingly precise 27, 31, 33, 34, 17, 19.
Fig.  1: Schematic representation of the problem under consideration. A spherical particle with radius $a$ and refractive index $n_{1}$ is embedded in an absorbing or gain medium with $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$, where $n_{2}^{\prime}$ and $n_{2}^{\prime\prime}$ are real. Imaginary refractive index of the host $n_{2}^{\prime\prime}>0$ for an absorbing medium and $n_{2}^{\prime\prime}<0$ for a gain medium, and $n_{2}^{\prime\prime}=0$ for a transparent medium. Normalized extinction coefficient $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$ is introduced to quantify the magnitude of the host dissipation or gain.
The outline of our contribution is as follows. Section 2 first recalls in its Subsection 2.1 the expression for distance-independent apparent extinction cross section $C_{\rm ext}$ in the framework of the Lorenz–Mie theory 13. In Subsection 2.2 our earlier developed MLWA 33, 34 is summarized. In Subsection 2.3, still within the framework of MLWA, an analytic first-order perturbation theory in a normalized extinction coefficient, $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$, representing the magnitude of the host dissipation, is developed. In Section 3, performance of the above approximations involving small plasmonic particles in different absorbing and active hosts is investigated. Discussion of some of the observed features is provided in Section 4. We then conclude with Section 5.
2 Theory
2.1 Apparent extinction cross section in the framework of the Lorenz–Mie theory
The specific expression for distance-independent apparent extinction cross section $C_{\rm ext}$ in the framework of the Lorenz–Mie theory is 13 is given as an infinite sum over different multipole orders $\ell\geq 1$,
$C_{\rm ext}=-\frac{2\pi}{k^{\prime}}\,\mbox{Re }\left\{\sum_{\ell=1}^{\infty}\frac{1}{k}\,(2\ell+1)(T_{E\ell}+T_{M\ell})\right\}.$ (1)
In the optical convention, $T_{p\ell}$ correspond to the familiar Mie’s expansion coefficients $a_{\ell}$ and $b_{\ell}$ (eqns (4.53) of ref. 2), i.e. $T_{E\ell}=-a_{\ell}$ for electric, or transverse magnetic (TM), polarization, and $T_{M\ell}=-b_{\ell}$ for magnetic, or transverse electric (TE), polarization. The physical, or operational, meaning of the distance-independent apparent $C_{\rm ext}$ is that it determines the reading of a polarization-sensitive well-collimated radiometer (WCR) at a sufficiently large distance $r$ from the particle 13,
$\mbox{WCR signal}\propto\exp(-2k^{\prime\prime}r)(\Omega-C_{\rm ext})I_{\rm inc},$
where $\Omega$ is the area of the objective lens of the WCR, and $I_{\rm inc}$ is the intensity of the incident homogeneous (uniform) plane wave at the center of the particle. Importantly, the distance-independent extinction cross-section cannot be introduced in the context of evaluating the energy budget of an arbitrarily shaped volume containing the scattering particle 13.
2.2 The dipole MLWA
Fig.  2: Demonstration of different physical mechanisms described within MLWA on the extinction spectra of Ag sphere with $a=40$ nm in the host with $n_{2}^{\prime}=1.33$. Here $\bar{\kappa}=0$ for a transparent host and $\bar{\kappa}=0.05$ in case of a dissipative host. The refractive indices of Ag were taken from ref. 35. (a) Quasi-static approximation; (b) radiative correction taking the effect of retardation with respect to the incident field; (c) the MLWA including additionally the dynamic depolarization term. Obviously all the correction terms of the dipole MLWA given by eqn (3) are necessary to achieve a fairly accurate approximation to the exact Mie theory shown in panel (d) for a comparison. All the above approximations, but the exact Mie theory, assume a constant field inside the sphere.
The MLWA is a rational approximation to the Mie coefficients in terms of a fraction of simple polynomials in size parameter $x$ that in a concise way combines three different elementary terms, involving size-independent quasi-static Fröhlich term,
$F_{p\ell}:=\upsilon+\frac{\ell+1}{\ell},$
the dynamic depolarization ($D_{p\ell}\sim x^{2}$) and the radiative reaction ($R_{p\ell}\sim x^{2\ell+1}$ for $p=E$ and $R_{p\ell}\sim x^{2\ell+3}$ for $p=M$) in the functional form 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19
$T_{p\ell}=\frac{iR_{p\ell}(x)}{F_{p\ell}+D_{p\ell}(x)-iR_{p\ell}(x)}\cdot$ (2)
Assuming nonmagnetic media, one has $\upsilon=\mu_{1}/\mu_{2}=1$ for magnetic (TE) polarization ($p=M$), and $\upsilon=\varepsilon:=\varepsilon_{1}/\varepsilon_{2}$ for electric (TM) polarization ($p=E$), where the subscript 1 (2) identifies the relevant quantities of a sphere (host). The functional form of eqn (2) makes it transparent that the usual Rayleigh limit, which amounts to setting $D_{p\ell}(x)=R_{p\ell}(x)\equiv 0$ in the denominator for $\ell=1$ and $p=E$, is recovered for $x,x_{s}\ll 1$. Here $x$ is in general complex size parameter, $x=2\pi a/\lambda$, with $\lambda$ being the wavelength in the host medium, whereas $x_{s}=2\pi n_{2}a/n_{1}\lambda=x/\sqrt{\varepsilon}$. The vanishing of the size-independent $F$ in the denominator yields the usual quasi-static Fröhlich LSPR condition, which determines the quasi-static LSPR frequencies $\omega_{0\ell}$. In what follows, we shall focus on the dipole MLWA which yields 33, 34
$T_{E1}\approx\frac{2i(\varepsilon-1)x^{3}/3}{\varepsilon+2-3(\varepsilon-2)x^{2}/5-2i(\varepsilon-1)x^{3}/3},$ (3)
where $\varepsilon=\varepsilon_{1}/\varepsilon_{2}$ is the relative dielectric function. An exceptional feature of the MLWA dipole contribution is that one can determine analytically an exact position of the complex pole of $T_{E1}$ (the Mie coefficient $-a_{1}$), and hence the dipolar LSPR position, at
$\varepsilon_{E1}=-2\times\frac{1+3x^{2}/5+ix^{3}/3}{1-3x^{2}/5-2ix^{3}/3},$ (4)
which corrects formula eqn (6) of ref. 19. The proof of that $\varepsilon_{E1}$ yields complex zero of the denominator $D$ of $a_{1}$ is relegated to Sec. S2 “Complex pole of the dipole contribution within the dipole MLWA” of ESI. In the limit of small $x$, one can expand the denominator of $\varepsilon_{E1}$ as $\sim 1+3x^{2}/5+2ix^{3}/3$, whereby eqn (4) reduces to the familiar classical Bohren and Huffman result (cf. eqn (12.13) of ref. 2):
$\varepsilon_{E1}\sim\varepsilon_{\rm BH}\approx-2-\frac{12x^{2}}{5}\qquad(|x|\ll 1).$ (5)
Fig.  3: Extinction spectra, $Q_{\rm ext}$, for individual spherical Ag nanoparticles with radii $a=10$, $20$, $30$, $40$, $50$ nm in water-like host ($n_{2}^{\prime}=1.33$) with $\bar{\kappa}=0.001,0.005,0.01,0.05$, $0.1$, as shown in the legend for each plot. Spectra were calculated by eqn (1) of the exact Mie theory (solid red line), the dipole MLWA by eqn (3) (blue dot-dashed line), the Mie theory in nonabsorbing host ($n_{2}^{\prime\prime}=0$; dashed green line), and the first order $\bar{\kappa}$-PT of our eqn (9) (dotted orange line). Because the Mie theory in the nonabsorbing host is independent of $\bar{\kappa}$, its plots are identical at each given column. The refractive indices of Ag were taken from ref. 35.
2.3 Perturbation theory in a normalized extinction coefficient within the MLWA
A useful parametrization of the relative dielectric function $\varepsilon$ between the sphere and the host, suitable for studying the departure from nonabsorbing host, is (cf. eqn (9) of ref. 19):
$\varepsilon=\frac{\varepsilon_{1}}{\varepsilon_{2}}=\frac{(n_{1}^{\prime}+in_{1}^{\prime\prime})^{2}}{(n_{2}^{\prime}+in_{2}^{\prime\prime})^{2}}=\frac{(n_{1}^{\prime}/n_{2}^{\prime}+in_{1}^{\prime\prime}/n_{2}^{\prime})^{2}}{(1+in_{2}^{\prime\prime}/n_{2}^{\prime})^{2}}=\frac{\varepsilon_{t}}{(1+i\bar{\kappa})^{2}},$ (6)
where $\varepsilon_{t}=(n_{1}^{\prime}/n_{2}^{\prime}+in_{1}^{\prime\prime}/n_{2}^{\prime})^{2}=\varepsilon_{1}/(n_{2}^{\prime})^{2}$ is the relative (in general complex number if $\varepsilon_{1}$ is complex) dielectric function between the sphere and a nonabsorbing host with real refractive index $n_{2}^{\prime}$, and $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$ is a normalized extinction coefficient representing the magnitude of the host dissipation. On substituting the Taylor expansion of the electric dipole term $a_{1}$,
$a_{1}(\varepsilon)=a_{1}(\varepsilon_{t})+\bar{\kappa}\left.\frac{{\rm d}a_{1}}{{\rm d}\bar{\kappa}}\right|_{\varepsilon=\varepsilon_{t}}+{\cal O}(\bar{\kappa}^{2}),$
in the expression eqn (1) of the apparent cross section, $C_{\rm ext}$, it is in principle possible to provide in a systematic way the results for the extinction efficiency, $Q_{\rm ext}=C_{\rm ext}/\pi a^{2}$, in the first order of $\bar{\kappa}$ for both gain and absorbing media in the dipole approximation. The derivative ${\rm d}a_{1}/{\rm d}\bar{\kappa}$ is determined by eqn (S21) of ESI (Section S5: “The derivative ${\rm d}a_{1}/{\rm d}\bar{\kappa}$”),
$\frac{{\rm d}a_{1}}{{\rm d}\bar{\kappa}}=-12n_{2}^{\prime}x^{\prime}\bar{\kappa}\,\mbox{Re}\left[\frac{2-\varepsilon_{t}(\varepsilon_{t}-1)+\frac{1}{5}(\varepsilon_{t}^{2}-\varepsilon_{t}+2)(x^{\prime})^{2}}{n_{2}D^{2}(\varepsilon_{t},x^{\prime})}\right],$ (7)
where $x^{\prime}=x_{0}n_{2}^{\prime}$, with $x_{0}=2\pi a/\lambda$ being the size parameter in vacuum host, and
$D(\varepsilon_{t},x^{\prime})=\varepsilon_{t}+2-3(\varepsilon_{t}-2)(x^{\prime})^{2}/5-2i(\varepsilon_{t}-1)(x^{\prime})^{3}/3$ (8)
is the denominator $D$ in eqn (3) in the limit $\bar{\kappa}\to 0$. One finds eventually
$Q_{\rm ext}$ $=$ $\frac{C_{\rm ext}}{\pi a^{2}}\approx\frac{2}{x^{\prime}}\,\mbox{Re}\left\{\frac{3}{x}\,\left[a_{1}(\varepsilon_{t})+\bar{\kappa}\left.\frac{da_{1}}{d\bar{\kappa}}\right|_{\varepsilon=\varepsilon_{t}}\right]\right\}$ (9)
$=$ $\frac{2}{x^{\prime}}\mbox{Re}\left\{\frac{3}{x}a_{1}(\varepsilon_{t})\right\}-$
[…]
which defines the first order expansion of $Q_{\rm ext}$ in the parameter $\bar{\kappa}$ within the perturbation theory (PT). The first term on the rhs is the dipole MLWA in a nonabsorbing host characterized by $\varepsilon_{t}$. The second term on the rhs is the perturbation correction. In what follows, we will refer to eqn (9) as the $\bar{\kappa}$-PT approximation.
3 Results
3.1 The dipole MLWA vs perturbation theory
Fig.  4: Extinction spectra, $Q_{\rm ext}$, for Al, Ag, Au and Mg spherical nanoparticles with $a=30$ nm embedded in a host absorption medium with $n_{2}^{\prime}=1.33$ (left column) and gain host medium (right column). The left (right) column is for $\bar{\kappa}=0.01$ ($\bar{\kappa}=-0.01$). $Q_{\rm ext}$ are shown as calculated by eqn (1) of the Mie theory (solid red line), within dipole MLWA of eqn (3) (blue dot-dashed line), the Mie theory approximation with the nonabsorbing host (green dashed line), and by the $\bar{\kappa}$-PT given by eqn (9) (dotted orange line). The refractive indices of Ag, Al, Au were taken from ref. 35 and those of Mg from ref. 36.
Fig. 3 compares performance of various approximations relative to the exact Mie theory involving
  • the dipole MLWA (with the sole term $T_{E1}$ in eqn (1) given by eqn (3))
  • the Mie theory approximation with nonabsorbing host (determined by eqn (1) with $n_{2}^{\prime}=0$)
  • the first-order $\bar{\kappa}$-PT given by eqn (9)
in water-like host ($n_{2}^{\prime}=1.33$) for different host absorption characterized by different values of $\bar{\kappa}$ on the example of Ag spheres with radii between $10$ and $50$ nm. Fig. 4 compares results for spheres made from most common plasmonic materials Al, Ag, Au, Mg with radius $a=30$ nm in water-like host ($n_{2}^{\prime}=1.33$) with $\bar{\kappa}=0.01$ and $-0.01$ (similar results for glass-like host ($n_{2}^{\prime}=1.5$) are presented in Fig. S2I of ESI). As demonstrated in those figures, the dipole MLWA can be very precise. In the case of Ag and Au nanoparticles, the dipole MLWA essentially overlies with the exact Mie theory results for $a\lesssim 25$ nm. The agreement, slightly better in the case of Au than Ag, continues to be acceptable up to $a\sim 50$ nm (Fig. S1III of ESI), and can be used, at least qualitatively, up to $a\sim 70$ nm (Fig. S1IV of ESI). In the case of Al and Mg nanoparticles, a slight deviation from Mie theory results becomes visible by the naked eye already for $a\gtrsim 25$ nm (cf. Figs. S1I and S1II of ESI). The agreement continues to be acceptable up to $a\sim 40$ nm and can be used at least qualitatively up to $a\sim 50$ nm (Fig. S1III of ESI).
A first general observation is that with increasing sphere radius the performance of the dipole MLWA worsens. This is not surprising, as the MLWA is by definition is a long wavelength approximation. The same applies to the $\bar{\kappa}$-PT. The $\bar{\kappa}$-PT overlies with the dipole MLWA for all $a\leq 50$ nm, provided that $\bar{\kappa}\leq 0.01$. The latter justifies a posteriori that our $\bar{\kappa}$-PT is correct. Another interesting tendency revealed by Fig. 3 (Figs. S1I and S1II of ESI) is that the maximal radius $a$ for which the dipole MLWA remains accurate increases with increasing $\bar{\kappa}$.
Fig.  5: Size dependence of $\bar{Q}_{\bar{\kappa}}$, the ratio of the PT correction to the leading term on the rhs of eqn (9), for Ag sphere at the (size-dependent) LSPR wavelength for different $\bar{\kappa}$: (a) $0.001$, (b) $0.005$, (c) $0.01$ and (d) $0.1$.
Fig.  6: Extinction spectra, $Q_{\rm ext}$, for spherical Ag nanoparticles with (a) $a=50$ nm, (b) $a=60$ nm, (c) $a=70$ nm, and (d) $a=80$ nm embedded in Poly(3-hexylthiophene) (P3HT) host medium. Spectra are shown as calculated with the exact Mie theory (solid red line), the Mie theory for non-absorbing media ($n_{2}^{\prime\prime}=0$) (green dashed line), the dipole MLWA (blue dot-dashed line), and the $\bar{\kappa}$-PT (dotted orange line).
Importantly, the results of $\bar{\kappa}$-PT and the dipole MLWA overlie with the exact Mie theory results for $a\lesssim 25$ nm, which supports earlier observation on the range of validity of the MLWA in an absorbing host by Khlebtsov 17. Whereas the dipole MLWA continues to overlie with the exact Mie theory results for $a\lesssim 25$ nm irrespective of $\bar{\kappa}$, the results of the $\bar{\kappa}$-PT begin to deviate from those of the dipole MLWA for $\bar{\kappa}\gtrsim 0.05$. This is to be expected, because the $\bar{\kappa}$-PT is a first order perturbation theory of the dipole MLWA in $\bar{\kappa}$.
Surprisingly, the smaller the particle, the greater the deviation of the $\bar{\kappa}$-PT relative to the dipole MLWA for $\bar{\kappa}\gtrsim 0.05$ (see Fig. S1 and S2 of ESI). The origin of this behaviour is that $D(\varepsilon_{t})$ is typically small in a proximity of a LSPR (cf. the generalized Fröhlich condition eqn (5)). Whereas the first term in eqn (9) is of the order $1/D(\varepsilon_{t})$, the term proportional to $\bar{\kappa}$ is of the order $1/D^{2}(\varepsilon_{t})$. However, when the size parameter $x$ increases, $|D(\varepsilon_{t})|$ at a LSPR decreases (due to a larger separation from the complex zero), and, as illustrated in Fig. 5, $\bar{Q}_{\bar{\kappa}}$, defined as the ratio of the PT correction to the leading term on the rhs of eqn (9), gradually decreases, whereby the first-order $\bar{\kappa}$-PT begins to approximate the MLWA. Note also how the ratio $\bar{Q}_{\bar{\kappa}}$ increases with $\bar{\kappa}$, which, as expected, explains why the $\bar{\kappa}$-PT begins to deviate from the MLWA with increasing $\bar{\kappa}$.
3.2 Mie theory approximation with nonabsorbing host
Similar behaviour is observed for the Mie theory approximation with nonabsorbing host ($n_{2}^{\prime\prime}=0$) that begins to deviate from the exact Mie theory results, but somewhat earlier, beginning with $\bar{\kappa}=0.01$. Again the smaller the particle, the greater the deviation. Nevertheless, the Mie theory approximation with nonabsorbing host becomes the best approximation with increasing particle radius (e.g. $a\gtrsim 40$ nm and $\bar{\kappa}\leq 0.01$). For example, the Mie theory in nonabsorbing host begins to overlie with the Mie theory in an absorbing host for $a\gtrsim 50$ nm and $\bar{\kappa}=0.01$ (see Fig. S1 of ESI). A threshold radius for which it happens increases with $\bar{\kappa}$.
Why the Mie theory approximation with nonabsorbing host becomes the best approximation for sufficiently large $a$ can be explained by increasing relevance of higher-order multipole contributions with increasing particle size. This is understandable, because with increasing $a$ higher order multipoles, which are obviously absent in any dipole approximation, become more and more relevant. This is demonstrated also in Fig. 6 showing the effect of increasing the radius of spherical Ag nanoparticles in Poly(3-hexylthiophene) (P3HT) host medium, which was used in a number of recent studies 15, 37, 17, 19. Unlike inherent dipole approximations, the Mie theory approximation with nonabsorbing host captures reasonably well both the dipole and quadrupole peaks. Obviously, higher-order MLWA of ref. 34 could have captured the quadrupole peak, but this goes beyond the scope of present study.
3.3 Isolating the host dissipative effects on the extinction cross sections
The dissipative effects of the host on the extinction cross sections, $Q_{\rm ext}$, can obviously be isolated by subtracting from the exact value $Q_{\rm ext}=Q_{{\rm ext};\varepsilon}$ the value of $Q_{{\rm ext};\varepsilon_{t}}$ obtained by the Mie theory for the non-absorbing host characterised by $n_{2}^{\prime\prime}=0$ (i.e. $\varepsilon\to\varepsilon_{t}$). In what follows, we denote the difference by $\eta_{\rm eff}:=Q_{{\rm ext};\varepsilon}-Q_{{\rm ext};\varepsilon_{t}}$. Thus $\eta_{\rm eff}$ quantifies the contribution of the host absorption or gain in the resulting $Q_{\rm ext}$. To our satisfaction, it turns out that, in a suitable parameter range, the analytic $\bar{\kappa}$-PT can reliably capture the effect of the host absorption on the extinction efficiency of a plasmonic nanosphere as demonstrated in Fig. 7. Not surprisingly, the agreement can also be reached in the case of gain media, which opens door for analysis of promising applications involving active media, see Fig. 4 (Fig. S5 of ESI) such as spasers 38, 39, 40. Therefore, within the range of its validity, the first order $\bar{\kappa}$-PT allows one to both intuitively and analytically understand the mechanisms of host dissipation or gain on the extinction efficiency of a plasmonic nanosphere.
Fig.  7: The effect $\eta_{\rm eff}$ of the host absorption (left column) and the gain host medium (right column) on $Q_{\rm ext}$ in the Mie theory (solid red line) and the $\bar{\kappa}$-PT (dashed dark green line) for Al, Ag, Au and Mg materials with different radii $a$ embedded in the host medium with $n_{2}^{\prime}=1.33$ and $\bar{\kappa}=0.001$ (left column) and $\bar{\kappa}=-0.001$ (right column).
4 Discussion
Rather exhaustive MLWA analysis in the absorbing case has been recently provided by Khlebtsov 17 by employing slightly different MLWA. The study of ref. 17 arrived at similar conclusions that MLWA is very precise for $a\lesssim 25$ nm. Our study complements ref. 17 with (i) a first order analytic perturbation theory around a nonabsorbing host in a normalized extinction coefficient $\bar{\kappa}$ ($\bar{\kappa}$-PT) and (ii) an investigation of gain media. In addition, we have examined the behaviour of Al and Mg nanoparticles.
5 Conclusions
Our previously developed modified dipole long-wave approximation (MLWA) was shown to essentially overly with the exact Mie theory results for $a\lesssim 25$ nm ($a\lesssim 20$ nm) in the case of Ag and Au (Al and Mg) nanoparticles. The agreement for Au and Ag (Al and Mg) nanoparticles, slightly better in the case of Au than Ag, continues to be acceptable up to $a\sim 50$ nm ($a\sim 40$ nm), and can be used, at least qualitatively, up to $a\sim 70$ nm ($a\sim 50$ nm). We developed within the dipole MLWA a first order analytic perturbation theory (PT) around a nonabsorbing host in a normalized extinction coefficient $\bar{\kappa}$ and investigated its properties. It was shown that, in a suitable parameter range, the $\bar{\kappa}$-PT can reliably capture the effect of host absorption or gain on the extinction efficiency of a plasmonic nanosphere. Considering growing interest in light-matter interactions, we expect that our results will help in designing optimal systems comprising plasmonic nanoparticles embedded in suitable dissipative or gain media for various applications.
Author Contributions
Conflicts of interest
There are no conflicts to declare.
Acknowledgements
Electronic Supplementary Information
The supplementary information includes Sec. S1 “Supplementary Figures”, where additional visual data are provided on apparent extinction efficiencies for further particle radii, host $n_{2}^{\prime}$, and normalized extinction coefficient $\bar{\kappa}$. In Sec. S2 “Complex Pole of the Dipole Contribution within the Dipole MLWA”, the predictive power of the complex pole knowledge on the LSPR position is investigated. Sec. S3 “Proof of that $\varepsilon_{E1}$ is the Complex Zero of the Denominator of $a_{1}$” presents mathematical proof of what is in its title. In Sec. S4 “The Imaginary Part of $\varepsilon_{E1}$ is Negative” proves the stated statement for any real $x$. In Sec. S5 “The Derivative $da_{1}/d\bar{\kappa}$” explicit calculations leading to eqn (7) are presented. Sec. S6 “A Previous Attempt to Formulate a Perturbation Theory (PT) in Terms of $\bar{\kappa}$” analyses an earlier failed theoretical attempt. Sec. S7 “Apparent Extinction Cross Section” reviews the absorption and scattering cross-sections within the MLWA. Finally, Sec. S8 “Conventional Scattering Theory” reviews some related aspects of conventional scattering theory.
Notes and references
  • Newton 1982 R. G. Newton, Scattering Theory of Waves and Particles, Springer Berlin Heidelberg, Berlin, Heidelberg, 1982.
  • Bohren and Huffman 1998 C. F. Bohren and D. R. Huffman, Absorption and Scattering of Light by Small Particles, Wiley-VCH Verlag GmbH & Co. KGaA, 1998.
  • Fuchs and Kliewer 1968 R. Fuchs and K. L. Kliewer, J. Opt. Soc. Am., 1968, 58, 319–330.
  • Mundy et al. 1974 W. C. Mundy, J. A. Roux and A. M. Smith, J. Opt. Soc. Am., 1974, 64, 1593–1597.
  • Chýlek and Pinnick 1979 P. Chýlek and R. G. Pinnick, Appl. Opt., 1979, 18, 1123–1124.
  • Bohren and Gilra 1979 C. F. Bohren and D. P. Gilra, J. Colloid Interface Sci., 1979, 72, 215–221.
  • Lebedev et al. 1999 A. N. Lebedev, M. Gartz, U. Kreibig and O. Stenzel, Eur. Phys. J. D., 1999, 6, 365–373.
  • Sudiarta and Chylek 2001 I. W. Sudiarta and P. Chylek, J. Opt. Soc. Am. A, 2001, 18, 1275–1278.
  • Yang et al. 2002 P. Yang, B.-C. Gao, W. J. Wiscombe, M. I. Mishchenko, S. E. Platnick, H.-L. Huang, B. A. Baum, Y. X. Hu, D. M. Winker, S.-C. Tsay and S. K. Park, Appl. Opt., 2002, 41, 2740–2759.
  • Fu and Sun 2006 Q. Fu and W. Sun, J. Quant. Spectrosc. Radiat. Transf., 2006, 100, 137–142.
  • Mishchenko 2007 M. I. Mishchenko, Opt. Express, 2007, 15, 13188–13201.
  • Mishchenko et al. 2017 M. I. Mishchenko, G. Videen and P. Yang, Opt. Lett., 2017, 42, 4873–4876.
  • Mishchenko and Yang 2018 M. I. Mishchenko and P. Yang, J. Quant. Spectrosc. Radiat. Transf., 2018, 205, 241–252.
  • Mishchenko and Dlugach 2019 M. I. Mishchenko and J. M. Dlugach, Appl. Opt., 2019, 58, 4871–4877.
  • Peck et al. 2019 R. L. Peck, A. G. Brolo and R. Gordon, J. Opt. Soc. Am. B, 2019, 36, F117–F122.
  • Mishchenko et al. 2019 M. I. Mishchenko, M. A. Yurkin and B. Cairns, OSA Continuum, 2019, 2, 2362–2368.
  • Khlebtsov 2021 N. G. Khlebtsov, Phys. Chem. Chem. Phys., 2021, 23, 23141–23157.
  • Dong et al. 2021 J. Dong, W. Zhang and L. Liu, Opt. Express, 2021, 29, 7690–7705.
  • Zhang et al. 2022 S. Zhang, J. Dong, W. Zhang, Minggang Luo and L. Liu, Opt. Lett., 2022, 47, 5577–5580.
  • Mishchenko and Dlugach 2019 M. I. Mishchenko and J. M. Dlugach, OSA Continuum, 2019, 2, 3415–3421.
  • Videen and Sun 2003 G. Videen and W. Sun, Appl. Opt., 2003, 42, 6724–6727.
  • Meier and Wokaun 1983 M. Meier and A. Wokaun, Opt. Lett., 1983, 8, 581–583.
  • Zeman and Schatz 1984 E. J. Zeman and G. C. Schatz, Dynamics on Surfaces, Dordrecht, 1984, pp. 413–424.
  • Zeman and Schatz 1987 E. J. Zeman and G. C. Schatz, J. Phys. Chem., 1987, 91, 634–643.
  • Kelly et al. 2003 K. L. Kelly, E. Coronado, L. L. Zhao and G. C. Schatz, J. Phys. Chem. B, 2003, 107, 668–677.
  • Kuwata et al. 2003 H. Kuwata, H. Tamaru, K. Esumi and K. Miyano, Appl. Phys. Lett., 2003, 83, 4625–4627.
  • Moroz 2009 A. Moroz, J. Opt. Soc. Am. B, 2009, 26, 517–527.
  • Zorić et al. 2011 I. Zorić, M. Zäch, B. Kasemo and C. Langhammer, ACS Nano, 2011, 5, 2535–2546.
  • Massa et al. 2013 E. Massa, S. A. Maier and V. Giannini, New J. Phys., 2013, 15, 063013.
  • Le Ru et al. 2013 E. C. Le Ru, W. R. C. Somerville and B. Auguié, Phys. Rev. A, 2013, 87, 012504.
  • Schebarchov et al. 2013 D. Schebarchov, B. Auguié and E. C. Le Ru, Phys. Chem. Chem. Phys., 2013, 15, 4233–4242.
  • Januar et al. 2020 M. Januar, B. Liu, J.-C. Cheng, K. Hatanaka, H. Misawa, H.-H. Hsiao and K.-C. Liu, J. Phys. Chem. C, 2020, 124, 3250–3259.
  • Rasskazov et al. 2020 I. L. Rasskazov, P. S. Carney and A. Moroz, Opt. Lett., 2020, 45, 4056–4059.
  • Rasskazov et al. 2021 I. L. Rasskazov, V. I. Zakomirnyi, A. D. Utyushev, P. S. Carney and A. Moroz, J. Phys. Chem. C, 2021, 125, 1963–1971.
  • McPeak et al. 2015 K. M. McPeak, S. V. Jayanti, S. J. P. Kress, S. Meyer, S. Iotti, A. Rossinelli and D. J. Norris, ACS Photonics, 2015, 2, 326–333.
  • Palm et al. 2018 K. J. Palm, J. B. Murray, T. C. Narayan and J. N. Munday, ACS Photonics, 2018, 5, 4677–4686.
  • Peck et al. 2021 R. Peck, A. Khademi, J. Ren, S. Hughes, A. G. Brolo and R. Gordon, Phys. Rev. Research, 2021, 3, 013014.
  • Bergman and Stockman 2003 D. J. Bergman and M. I. Stockman, Phys. Rev. Lett., 2003, 90, 027402.
  • Lawandy 2004 N. M. Lawandy, Appl. Phys. Lett., 2004, 85, 5040–5042.
  • Veltri et al. 2016 A. Veltri, A. Chipouline and A. Aradian, Sci. Rep., 2016, 6, 33018.

### v2

@@
-Anton D. Utyushev,a Vadim I. Zakomirnyi,b Alexey A. Shcherbakov,a Ilia L. Rasskazovc and Alexander Morozd
+Intuitive understanding of extinction of small particles in absorbing and active host media within the MLWA
+Intuitive understanding of extinction of small particles in absorbing and active host media within the MLWA
 […]
-[Uncaptioned image] Intuitive understanding of extinction of small particles in absorbing and active host media within the MLWA
-Anton D. Utyushev,a Vadim I. Zakomirnyi,b Alexey A. Shcherbakov,a Ilia L. Rasskazovc and Alexander Morozd
-[Uncaptioned image] In an absorbing or an active host medium characterized by a complex refractive index $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$, our previously developed modified dipole long-wave approximation (MLWA) is shown to essentially overly with the exact Mie theory results for spherical nanoparticle with radius $a\lesssim 25$ nm ($a\lesssim 20$ nm) in the case of Ag and Au (Al and Mg) nanoparticles.
+[…]
+[…]
+[…]
+[…]
+[…]
+journal: opticajournalarticletype: Research Article
+[…]
+{abstract*}
+In an absorbing or an active host medium characterized by a complex refractive index $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$, our previously developed modified dipole long-wave approximation (MLWA) is shown to essentially overly with the exact Mie theory results for localized surface plasmon resonance of spherical nanoparticles with radius $a\lesssim 25$ nm ($a\lesssim 20$ nm) in the case of Ag and Au (Al and Mg) nanoparticles.
 The agreement for Au and Ag (Al and Mg) nanoparticles, slightly better in the case of Au than Ag, continues to be acceptable up to $a\sim 50$ nm ($a\sim 40$ nm), and can be used, at least qualitatively, up to $a\sim 70$ nm ($a\sim 50$ nm) correspondingly.
 A first order analytic perturbation theory (PT) in a normalized extinction coefficient, $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$, around a nonabsorbing host is developed within the dipole MLWA and its properties are investigated.
-It is shown that, in a suitable parameter range, the PT can reliably capture the effect of host absorption or gain on the extinction efficiency of various plasmonic nanoparticles.
-footnotetext: a School of Physics and Engineering, ITMO University, 197101, Saint-Petersburg, Russiafootnotetext: b Beckman Institute for Advanced Science and Technology, University of Illinois at Urbana-Champaign, Urbana, IL 61801, USAfootnotetext: c SunDensity Inc., Rochester, NY 14604, USA footnotetext: d Wave-scattering.com footnotetext:  Corresponding author: ad.utyushev@gmail.comfootnotetext: † Electronic Supplementary Information (ESI) available: See DOI:
+It is shown that, in a suitable parameter range, the PT can reliably isolate and capture the effect of host absorption or host gain on the overall extinction efficiency of various plasmonic nanoparticles.
 1 Introduction
 Electromagnetic scattering in an absorbing host characterized by a complex refractive index $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$ (where $n_{2}^{\prime}$ and $n_{2}^{\prime\prime}$ are real) has more than fifty years old history.
-The traditional scattering theory neglects the host dissipation and gain 1, 2, because those cases imply either vanishing or infinite scattering wave at the spatial infinity.
-Once wave number $k=k^{\prime}+{\rm i}k^{\prime\prime}$ is a complex number, conventional expressions for cross sections cannot be straightforwardly extended for $k^{\prime\prime}\neq 0$, because the expressions yield cross sections as complex quantities.
-Not surprisingly, the history of scattering in an absorbing host is filled in with a number of controversies 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19.
+The traditional scattering theory neglects the host dissipation and gain [1, 2], because those cases imply either vanishing or infinite scattering wave at the spatial infinity.
+Once wave number $k=k^{\prime}+{\rm i}k^{\prime\prime}=2\pi n_{2}/\lambda_{0}$, where $\lambda_{0}$ is the vacuum wavelength, is a complex number, i.e.
+$k^{\prime\prime}\neq 0$, conventional expressions for cross sections cannot be straightforwardly extended for $k^{\prime\prime}\neq 0$, because the expressions yield cross sections as complex quantities.
+Not surprisingly, the history of scattering in an absorbing host is filled in with a number of controversies [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
 This is probably why in classical textbooks it is only fleetingly mentioned in Section 12.1.3 of ref.
 2.
-Already the definition of an incident intensity is not straightforward, as the field incident on a particle is different at different points on the particle as a result of an absorbing medium 4, 5.
-In an absorbing or an active unbounded host, the imaginary part $k^{\prime\prime}$ of the complex wave number $k=k^{\prime}+ik^{\prime\prime}=2\pi n_{2}/\lambda_{0}$, where $\lambda_{0}$ is the vacuum wavelength, is nonzero, $k^{\prime\prime}\neq 0$.
-A great deal of effort was required to arrive at suitable definitions of cross sections for $k^{\prime\prime}\neq 0$.
+Already the definition of an incident intensity is not straightforward, as the field incident on a particle is different at different points on the particle as a result of an absorbing medium [4, 5].
 In contrast to the conventional case of $k^{\prime\prime}=0$, two sets of cross sections are commonly used: inherent and apparent.
 The former is obtained by performing surface integrals of corresponding Poynting vectors over the particle surface.
-The approach was developed a long time ago 4, 5, but currently accepted expressions for the inherent cross-sections were not presented until 1999 7, 8, 9, 10.
-The history of apparent cross sections, which are operationally defined in the far field, began with the optical theorem of Bohren and Gilra 6.
-Here, too, it took a long time, until 2007, to arrive at definitive expressions for the apparent extinction cross-section cross section, $C_{\rm ext}$, and the so-called apparent scattering cross section, $C_{\rm sca}$ 11, 12, 13, 14.
-A curious feature of the apparent extinction cross-section in an absorbing host is that it can be negative 20, which is neither an artifact of numerical simulations nor it violates any physical law.
+The approach was developed a long time ago [4, 5], but currently accepted expressions for the inherent cross-sections were not presented until 1999 [7, 8, 9, 10].
+The focus of present work will be on the apparent extinction cross-section $C_{\rm ext}$, which is operationally defined in the far field (see Section 2.1 below).
+The history of apparent cross sections began with the optical theorem of Bohren and Gilra [6].
+Here, too, it took a long time, until 2007, to arrive at definitive expressions for the apparent extinction cross-section cross section, $C_{\rm ext}$ [11, 12, 13, 14].
+A curious feature of the apparent extinction cross-section in an absorbing host is that it can be negative [20], which is neither an artifact of numerical simulations nor it violates any physical law.
 The apparent extinction cross-section quantifies the difference in the readings of a forward-scattering detector taken with and without the particle.
 If the surrounding medium is absorbing, the presence of the particle can in fact make the detector signal stronger, thereby implying a negative extinction cross-section.
-There is no violation of the energy conservation law, since in this case the extinction cross-section is not used to quantify the energy budget of a finite volume encompassing the particle 20.
+There is no violation of the energy conservation law, since in this case the extinction cross-section is not used to quantify the energy budget of a finite volume encompassing the particle [20].
 Another curiosity of apparent cross-sections in an absorbing host is that an intrinsic definition of an apparent absorption cross-sections, $C_{\rm abs}$, is still missing.
 The difficulty lies in that the very presence of a particle necessarily modifies the near field around the particle.
-The latter may be the cause of an additional absorption in the host medium outside the particle compared to what is happening in the absence of the particle 8, 9.
-The critical point is that because the field is disturbed by the particle, additional absorption may be realized in the medium external to the particle 21.
-First-principles far-field computations based on the general Lorenz–Mie theory showed that increasing absorption in the host medium broadens and suppresses plasmon resonances in the apparent extinction and effective scattering efficiency factors and suppresses resonance features in the phase function 14.
-At the same time, using inherent cross-sections it was shown that a surrounding lossy medium can narrow the plasmonic resonances, as well as increase their amplitude dramatically 15, 17.
-Given this recent revival of interest in the role of the host dissipation or gain on extinction cross sections with conflicting predictions of it either enhancing or reducing the localized surface plasmonic resonance (LSPR) of small particles 14, 15, 17, 18, 19, the focus of present work is on the apparent extinction cross-section $C_{\rm ext}$ in the so-called modified long-wave approximation (MLWA) 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19.
-The MLWA 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19, which can be viewed as a next-order approximation beyond the Rayleigh limit, is known to overcome a number of severe deficiencies of the Rayleigh limit (see eqn (2) below) and, at least for nonabsorbing hosts, be surprisingly precise 27, 31, 33, 34, 17, 19.
-Fig.
-1: Schematic representation of the problem under consideration.
+The latter may be the cause of an additional absorption in the host medium outside the particle compared to what is happening in the absence of the particle [8, 9].
+The critical point is that because the field is disturbed by the particle, additional absorption may be realized in the medium external to the particle [21].
+Consequently, unlike $C_{\rm abs}$ and $C_{\rm sca}$, only apparent extinction cross-section $C_{\rm ext}$ can be consistently defined (see Section 2.1 below).
+The focus of present work is on the apparent extinction cross-section $C_{\rm ext}$ in the so-called modified long-wave approximation (MLWA) [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19].
+Firstly, it will be shown that the apparent extinction cross-section conforms better to physical intuition than the intrinsic extinction cross-section.
+Secondly, the MLWA [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19], which can be viewed as a next-order approximation beyond the Rayleigh limit, is known to overcome a number of severe deficiencies of the Rayleigh limit (see Eq.
+(2) below) and, at least for nonabsorbing hosts, be surprisingly precise [27, 31, 33, 34, 17, 19].
+Figure 1: Schematic representation of the problem under consideration.
 A spherical particle with radius $a$ and refractive index $n_{1}$ is embedded in an absorbing or gain medium with $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$, where $n_{2}^{\prime}$ and $n_{2}^{\prime\prime}$ are real.
@@
 The outline of our contribution is as follows.
-Section 2 first recalls in its Subsection 2.1 the expression for distance-independent apparent extinction cross section $C_{\rm ext}$ in the framework of the Lorenz–Mie theory 13.
-In Subsection 2.2 our earlier developed MLWA 33, 34 is summarized.
-In Subsection 2.3, still within the framework of MLWA, an analytic first-order perturbation theory in a normalized extinction coefficient, $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$, representing the magnitude of the host dissipation, is developed.
-In Section 3, performance of the above approximations involving small plasmonic particles in different absorbing and active hosts is investigated.
+Section 2 first recalls in its Subsection 2.1 the expression for distance-independent apparent extinction cross section $C_{\rm ext}$ in the framework of the Lorenz–Mie theory [13].
+In Subsection 2.2 our earlier developed MLWA [33, 34] is summarized.
+In Subsection 2.3, still within the framework of MLWA, an analytic first-order perturbation theory (PT) in a normalized extinction coefficient, $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$, representing the magnitude of the host dissipation, is developed.
+A motivation for developing the above PT is to isolate and capture the effect of the host absorption or host gain on the overall extinction, which is not possible within the MLWA (the latter captures only overall extinction).
+In Section 3, performance of the above approximations in describing localized surface plasmon resonances of spherical plasmonic nanoparticles in different absorbing and active hosts is investigated.
 Discussion of some of the observed features is provided in Section 4.
@@
 2.1 Apparent extinction cross section in the framework of the Lorenz–Mie theory
-The specific expression for distance-independent apparent extinction cross section $C_{\rm ext}$ in the framework of the Lorenz–Mie theory is 13 is given as an infinite sum over different multipole orders $\ell\geq 1$,
+The specific expression for distance-independent apparent extinction cross section $C_{\rm ext}$ in the framework of the Lorenz–Mie theory is [13] is given as an infinite sum over different multipole orders $\ell\geq 1$ (see Fig.
+1),
 $C_{\rm ext}=-\frac{2\pi}{k^{\prime}}\,\mbox{Re }\left\{\sum_{\ell=1}^{\infty}\frac{1}{k}\,(2\ell+1)(T_{E\ell}+T_{M\ell})\right\}.$ (1)
-In the optical convention, $T_{p\ell}$ correspond to the familiar Mie’s expansion coefficients $a_{\ell}$ and $b_{\ell}$ (eqns (4.53) of ref.
+In the optical convention, $T_{p\ell}$ correspond to the familiar Mie’s expansion coefficients $a_{\ell}$ and $b_{\ell}$ (Eqs.
+(4.53) of ref.
 2), i.e.
 $T_{E\ell}=-a_{\ell}$ for electric, or transverse magnetic (TM), polarization, and $T_{M\ell}=-b_{\ell}$ for magnetic, or transverse electric (TE), polarization.
-The physical, or operational, meaning of the distance-independent apparent $C_{\rm ext}$ is that it determines the reading of a polarization-sensitive well-collimated radiometer (WCR) at a sufficiently large distance $r$ from the particle 13,
+The physical, or operational, meaning of the distance-independent apparent $C_{\rm ext}$ is that it determines the reading of a polarization-sensitive well-collimated radiometer (WCR) at a sufficiently large distance $r$ from the particle [13],
 $\mbox{WCR signal}\propto\exp(-2k^{\prime\prime}r)(\Omega-C_{\rm ext})I_{\rm inc},$
 where $\Omega$ is the area of the objective lens of the WCR, and $I_{\rm inc}$ is the intensity of the incident homogeneous (uniform) plane wave at the center of the particle.
-Importantly, the distance-independent extinction cross-section cannot be introduced in the context of evaluating the energy budget of an arbitrarily shaped volume containing the scattering particle 13.
+Importantly, the distance-independent extinction cross-section cannot be introduced in the context of evaluating the energy budget of an arbitrarily shaped volume containing the scattering particle [13].
 2.2 The dipole MLWA
-Fig.
-2: Demonstration of different physical mechanisms described within MLWA on the extinction spectra of Ag sphere with $a=40$ nm in the host with $n_{2}^{\prime}=1.33$.
+Figure 2: Demonstration of different physical mechanisms described within MLWA on the extinction spectra of Ag sphere with $a=40$ nm in the host with $n_{2}^{\prime}=1.33$.
 Here $\bar{\kappa}=0$ for a transparent host and $\bar{\kappa}=0.05$ in case of a dissipative host.
@@
 35.
-(a) Quasi-static approximation; (b) radiative correction taking the effect of retardation with respect to the incident field; (c) the MLWA including additionally the dynamic depolarization term.
-Obviously all the correction terms of the dipole MLWA given by eqn (3) are necessary to achieve a fairly accurate approximation to the exact Mie theory shown in panel (d) for a comparison.
+(a) Quasi-static Rayleigh approximation; (b) radiative correction taking the effect of retardation with respect to the incident field; (c) the MLWA including additionally the dynamic depolarization term.
+Obviously all the correction terms of the dipole MLWA given by Eq.
+(3) are necessary to achieve a fairly accurate approximation to the exact Mie theory shown in panel (d) for a comparison.
 All the above approximations, but the exact Mie theory, assume a constant field inside the sphere.
@@
 $F_{p\ell}:=\upsilon+\frac{\ell+1}{\ell},$
-the dynamic depolarization ($D_{p\ell}\sim x^{2}$) and the radiative reaction ($R_{p\ell}\sim x^{2\ell+1}$ for $p=E$ and $R_{p\ell}\sim x^{2\ell+3}$ for $p=M$) in the functional form 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19
+the dynamic depolarization ($D_{p\ell}\sim x^{2}$) and the radiative reaction ($R_{p\ell}\sim x^{2\ell+1}$ for $p=E$ and $R_{p\ell}\sim x^{2\ell+3}$ for $p=M$) in the functional form [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19]
 $T_{p\ell}=\frac{iR_{p\ell}(x)}{F_{p\ell}+D_{p\ell}(x)-iR_{p\ell}(x)}\cdot$ (2)
 Assuming nonmagnetic media, one has $\upsilon=\mu_{1}/\mu_{2}=1$ for magnetic (TE) polarization ($p=M$), and $\upsilon=\varepsilon:=\varepsilon_{1}/\varepsilon_{2}$ for electric (TM) polarization ($p=E$), where the subscript 1 (2) identifies the relevant quantities of a sphere (host).
-The functional form of eqn (2) makes it transparent that the usual Rayleigh limit, which amounts to setting $D_{p\ell}(x)=R_{p\ell}(x)\equiv 0$ in the denominator for $\ell=1$ and $p=E$, is recovered for $x,x_{s}\ll 1$.
+The respective terms $F_{p\ell}$, $D_{p\ell}$, $R_{p\ell}$ in the functional form of Eq.
+(2) have well-defined physical origin and meaning (see Fig.
+2).
+This allows for an intuitive understanding of scattering from small particles which, as it will be shown, allows for rather reliable substitution of infinite series expansion in terms of Bessel functions of the conventional Mie solution by a single dipole term.
+The functional form of Eq.
+(2) makes it transparent that the usual Rayleigh limit, which amounts to setting $D_{p\ell}(x)=R_{p\ell}(x)\equiv 0$ in the denominator for $\ell=1$ and $p=E$, is recovered for $x,x_{s}\ll 1$.
 Here $x$ is in general complex size parameter, $x=2\pi a/\lambda$, with $\lambda$ being the wavelength in the host medium, whereas $x_{s}=2\pi n_{2}a/n_{1}\lambda=x/\sqrt{\varepsilon}$.
-The vanishing of the size-independent $F$ in the denominator yields the usual quasi-static Fröhlich LSPR condition, which determines the quasi-static LSPR frequencies $\omega_{0\ell}$.
-In what follows, we shall focus on the dipole MLWA which yields 33, 34
+The vanishing of the size-independent $F$ in the denominator yields the usual quasi-static Fröhlich condition which determines the quasi-static frequencies $\omega_{0\ell}$ for the occurrence of a localized surface plasmon resonance (LSPR).
+In what follows, we shall focus on the dipole MLWA which yields [33, 34]
 $T_{E1}\approx\frac{2i(\varepsilon-1)x^{3}/3}{\varepsilon+2-3(\varepsilon-2)x^{2}/5-2i(\varepsilon-1)x^{3}/3},$ (3)
@@
 $\varepsilon_{E1}=-2\times\frac{1+3x^{2}/5+ix^{3}/3}{1-3x^{2}/5-2ix^{3}/3},$ (4)
-which corrects formula eqn (6) of ref.
+which corrects formula Eq.
+(6) of ref.
 19.
-The proof of that $\varepsilon_{E1}$ yields complex zero of the denominator $D$ of $a_{1}$ is relegated to Sec.
-S2 “Complex pole of the dipole contribution within the dipole MLWA” of ESI.
-In the limit of small $x$, one can expand the denominator of $\varepsilon_{E1}$ as $\sim 1+3x^{2}/5+2ix^{3}/3$, whereby eqn (4) reduces to the familiar classical Bohren and Huffman result (cf.
-eqn (12.13) of ref.
+The proof of that $\varepsilon_{E1}$ yields complex zero of the denominator $D$ of $a_{1}$ is relegated to Section S3 in Supplement 1.
+In the limit of small $x$, one can expand the denominator of $\varepsilon_{E1}$ as $\sim 1+3x^{2}/5+2ix^{3}/3$, whereby Eq.
+(4) reduces to the familiar classical Bohren and Huffman result (cf.
+Eq.
+(12.13) of ref.
 2):
 $\varepsilon_{E1}\sim\varepsilon_{\rm BH}\approx-2-\frac{12x^{2}}{5}\qquad(|x|\ll 1).$ (5)
-Fig.
-3: Extinction spectra, $Q_{\rm ext}$, for individual spherical Ag nanoparticles with radii $a=10$, $20$, $30$, $40$, $50$ nm in water-like host ($n_{2}^{\prime}=1.33$) with $\bar{\kappa}=0.001,0.005,0.01,0.05$, $0.1$, as shown in the legend for each plot.
-Spectra were calculated by eqn (1) of the exact Mie theory (solid red line), the dipole MLWA by eqn (3) (blue dot-dashed line), the Mie theory in nonabsorbing host ($n_{2}^{\prime\prime}=0$; dashed green line), and the first order $\bar{\kappa}$-PT of our eqn (9) (dotted orange line).
+Figure 3: Extinction spectra, $Q_{\rm ext}$, for individual spherical Ag nanoparticles with radii $a=10$, $20$, $30$, $40$, $50$ nm in water-like host ($n_{2}^{\prime}=1.33$) with $\bar{\kappa}=0.001$, $0.005$, $0.01$, $0.05$, $0.1$, as shown in the legend for each plot.
+Spectra were calculated by Eq.
+(1) of the exact Mie theory (solid red line), the dipole MLWA by Eq.
+(3) (blue dot-dashed line), the Mie theory in nonabsorbing host ($n_{2}^{\prime\prime}=0$; dashed green line), and the first order $\bar{\kappa}$-PT of our Eq.
+(9) (dotted orange line).
 Because the Mie theory in the nonabsorbing host is independent of $\bar{\kappa}$, its plots are identical at each given column.
@@
 35.
+Note a broadening and suppression of plasmon resonances with increasing host absorption.
 2.3 Perturbation theory in a normalized extinction coefficient within the MLWA
-A useful parametrization of the relative dielectric function $\varepsilon$ between the sphere and the host, suitable for studying the departure from nonabsorbing host, is (cf. eqn (9) of ref. 19):
+The motivation for developing the perturbation theory is to isolate and capture the effect of the host absorption or host gain on the overall extinction, which is not possible within the MLWA (the latter captures only overall extinction).
+To this end, a useful parametrization of the relative dielectric function $\varepsilon$ between the sphere and the host, suitable for studying the departure from nonabsorbing host, is (cf.
+Eq.
+(9) of ref.
+19):
 $\varepsilon=\frac{\varepsilon_{1}}{\varepsilon_{2}}=\frac{(n_{1}^{\prime}+in_{1}^{\prime\prime})^{2}}{(n_{2}^{\prime}+in_{2}^{\prime\prime})^{2}}=\frac{(n_{1}^{\prime}/n_{2}^{\prime}+in_{1}^{\prime\prime}/n_{2}^{\prime})^{2}}{(1+in_{2}^{\prime\prime}/n_{2}^{\prime})^{2}}=\frac{\varepsilon_{t}}{(1+i\bar{\kappa})^{2}},$ (6)
@@
 $a_{1}(\varepsilon)=a_{1}(\varepsilon_{t})+\bar{\kappa}\left.\frac{{\rm d}a_{1}}{{\rm d}\bar{\kappa}}\right|_{\varepsilon=\varepsilon_{t}}+{\cal O}(\bar{\kappa}^{2}),$
-in the expression eqn (1) of the apparent cross section, $C_{\rm ext}$, it is in principle possible to provide in a systematic way the results for the extinction efficiency, $Q_{\rm ext}=C_{\rm ext}/\pi a^{2}$, in the first order of $\bar{\kappa}$ for both gain and absorbing media in the dipole approximation.
-The derivative ${\rm d}a_{1}/{\rm d}\bar{\kappa}$ is determined by eqn (S21) of ESI (Section S5: “The derivative ${\rm d}a_{1}/{\rm d}\bar{\kappa}$”),
+in the expression Eq.
+(1) of the apparent cross section, $C_{\rm ext}$, it is in principle possible to provide in a systematic way the results for the extinction efficiency, $Q_{\rm ext}=C_{\rm ext}/\pi a^{2}$, in the first order of $\bar{\kappa}$ for both gain and absorbing media in the dipole approximation.
+The derivative ${\rm d}a_{1}/{\rm d}\bar{\kappa}$ is determined by Eq.
+S21 from Section S5 in Supplement 1,
 $\frac{{\rm d}a_{1}}{{\rm d}\bar{\kappa}}=-12n_{2}^{\prime}x^{\prime}\bar{\kappa}\,\mbox{Re}\left[\frac{2-\varepsilon_{t}(\varepsilon_{t}-1)+\frac{1}{5}(\varepsilon_{t}^{2}-\varepsilon_{t}+2)(x^{\prime})^{2}}{n_{2}D^{2}(\varepsilon_{t},x^{\prime})}\right],$ (7)
@@
 $D(\varepsilon_{t},x^{\prime})=\varepsilon_{t}+2-3(\varepsilon_{t}-2)(x^{\prime})^{2}/5-2i(\varepsilon_{t}-1)(x^{\prime})^{3}/3$ (8)
-is the denominator $D$ in eqn (3) in the limit $\bar{\kappa}\to 0$. One finds eventually
+is the denominator $D$ in Eq. (3) in the limit $\bar{\kappa}\to 0$. One finds eventually
 $Q_{\rm ext}$ $=$ $\frac{C_{\rm ext}}{\pi a^{2}}\approx\frac{2}{x^{\prime}}\,\mbox{Re}\left\{\frac{3}{x}\,\left[a_{1}(\varepsilon_{t})+\bar{\kappa}\left.\frac{da_{1}}{d\bar{\kappa}}\right|_{\varepsilon=\varepsilon_{t}}\right]\right\}$ (9)
@@
 The second term on the rhs is the perturbation correction.
-In what follows, we will refer to eqn (9) as the $\bar{\kappa}$-PT approximation.
+In what follows, we will refer to Eq.
+(9) as the $\bar{\kappa}$-PT approximation.
 3 Results
 3.1 The dipole MLWA vs perturbation theory
-Fig.
-4: Extinction spectra, $Q_{\rm ext}$, for Al, Ag, Au and Mg spherical nanoparticles with $a=30$ nm embedded in a host absorption medium with $n_{2}^{\prime}=1.33$ (left column) and gain host medium (right column).
-The left (right) column is for $\bar{\kappa}=0.01$ ($\bar{\kappa}=-0.01$).
-$Q_{\rm ext}$ are shown as calculated by eqn (1) of the Mie theory (solid red line), within dipole MLWA of eqn (3) (blue dot-dashed line), the Mie theory approximation with the nonabsorbing host (green dashed line), and by the $\bar{\kappa}$-PT given by eqn (9) (dotted orange line).
+Figure 4: Extinction spectra, $Q_{\rm ext}$, for Al, Ag, Au and Mg spherical nanoparticles with $a=25$ nm embedded in a host absorption medium with $n_{2}^{\prime}=1.33$ (left two columns) and gain host medium (right two columns) for $\bar{\kappa}=0.01,0.1$ ($\bar{\kappa}=-0.01,-0.1$).
+$Q_{\rm ext}$ are shown as calculated by Eq.
+(1) of the Mie theory (solid red line), within dipole MLWA of Eq.
+(3) (blue dot-dashed line), the Mie theory approximation with the nonabsorbing host (green dashed line), and by the $\bar{\kappa}$-PT given by Eq.
+(9) (dotted orange line).
 The refractive indices of Ag, Al, Au were taken from ref.
@@
 36.
+Whereas absorption decreases the height of a LSPR and broadens its width, the gain does just the opposite.
+Note enlarged scale on the vertical axis in the last column for $\bar{\kappa}=-0.1$.
 Fig. 3 compares performance of various approximations relative to the exact Mie theory involving
-  • the dipole MLWA (with the sole term $T_{E1}$ in eqn (1) given by eqn (3))
-  • the Mie theory approximation with nonabsorbing host (determined by eqn (1) with $n_{2}^{\prime}=0$)
-  • the first-order $\bar{\kappa}$-PT given by eqn (9)
+  • the dipole MLWA (with the sole term $T_{E1}$ in Eq. (1) given by Eq. (3))
+  • the Mie theory approximation with nonabsorbing host (determined by Eq. (1) with $n_{2}^{\prime\prime}=0$)
+  • the first-order $\bar{\kappa}$-PT given by Eq. (9)
 in water-like host ($n_{2}^{\prime}=1.33$) for different host absorption characterized by different values of $\bar{\kappa}$ on the example of Ag spheres with radii between $10$ and $50$ nm.
 Fig.
-4 compares results for spheres made from most common plasmonic materials Al, Ag, Au, Mg with radius $a=30$ nm in water-like host ($n_{2}^{\prime}=1.33$) with $\bar{\kappa}=0.01$ and $-0.01$ (similar results for glass-like host ($n_{2}^{\prime}=1.5$) are presented in Fig.
-S2I of ESI).
+4 compares results for spheres made from most common plasmonic materials Al, Ag, Au, Mg with radius $a=25$ nm in water-like host ($n_{2}^{\prime}=1.33$) with $\bar{\kappa}=\pm 0.01$ and ±$\pm 0.1$ (similar results for glass-like host ($n_{2}^{\prime}=1.5$) are presented in Fig.
+S2 in Supplement 1).
+For different metals of plasmonic particles tabulated data of refractive indices have been used as follows: Al [35], Ag [35], Au [35], Mg [36].
 As demonstrated in those figures, the dipole MLWA can be very precise.
@@
 The agreement, slightly better in the case of Au than Ag, continues to be acceptable up to $a\sim 50$ nm (Fig.
-S1III of ESI), and can be used, at least qualitatively, up to $a\sim 70$ nm (Fig.
-S1IV of ESI).
+S1III in Supplement 1), and can be used, at least qualitatively, up to $a\sim 70$ nm (Fig.
+S1IV in Supplement 1).
 In the case of Al and Mg nanoparticles, a slight deviation from Mie theory results becomes visible by the naked eye already for $a\gtrsim 25$ nm (cf.
 Figs.
-S1I and S1II of ESI).
+S1I and S1II in Supplement 1).
 The agreement continues to be acceptable up to $a\sim 40$ nm and can be used at least qualitatively up to $a\sim 50$ nm (Fig.
-S1III of ESI).
+S1III in Supplement 1).
 A first general observation is that with increasing sphere radius the performance of the dipole MLWA worsens.
@@
 3 (Figs.
-S1I and S1II of ESI) is that the maximal radius $a$ for which the dipole MLWA remains accurate increases with increasing $\bar{\kappa}$.
-Fig.
-5: Size dependence of $\bar{Q}_{\bar{\kappa}}$, the ratio of the PT correction to the leading term on the rhs of eqn (9), for Ag sphere at the (size-dependent) LSPR wavelength for different $\bar{\kappa}$: (a) $0.001$, (b) $0.005$, (c) $0.01$ and (d) $0.1$.
-Fig.
-6: Extinction spectra, $Q_{\rm ext}$, for spherical Ag nanoparticles with (a) $a=50$ nm, (b) $a=60$ nm, (c) $a=70$ nm, and (d) $a=80$ nm embedded in Poly(3-hexylthiophene) (P3HT) host medium.
+S1I and S1II in Supplement 1) is that the maximal radius $a$ for which the dipole MLWA remains accurate increases with increasing $\bar{\kappa}$.
+This observation can be explained by the fact that for a given particle radius an increasing host absorption suppresses the contribution of higher-order multipoles that would normally arise at larger particle sizes.
+This allows the dipole MLWA to remain valid over a wider range of particle sizes, as the multipole effects are suppressed in absorbing media.
+Figure 5: Size dependence of $\bar{Q}_{\bar{\kappa}}$, the ratio of the PT correction to the leading term on the rhs of Eq.
+(9), for Ag sphere at the (size-dependent) LSPR wavelength for different $\bar{\kappa}$: (a) $0.001$, (b) $0.005$, (c) $0.01$ and (d) $0.1$.
+Figure 6: Extinction spectra, $Q_{\rm ext}$, for spherical Ag nanoparticles with (a) $a=50$ nm, (b) $a=60$ nm, (c) $a=70$ nm, and (d) $a=80$ nm embedded in Poly(3-hexylthiophene) (P3HT) host medium.
 Spectra are shown as calculated with the exact Mie theory (solid red line), the Mie theory for non-absorbing media ($n_{2}^{\prime\prime}=0$) (green dashed line), the dipole MLWA (blue dot-dashed line), and the $\bar{\kappa}$-PT (dotted orange line).
-Importantly, the results of $\bar{\kappa}$-PT and the dipole MLWA overlie with the exact Mie theory results for $a\lesssim 25$ nm, which supports earlier observation on the range of validity of the MLWA in an absorbing host by Khlebtsov 17.
+Importantly, the results of $\bar{\kappa}$-PT and the dipole MLWA overlie with the exact Mie theory results for $a\lesssim 25$ nm, which supports earlier observation on the range of validity of the MLWA in an absorbing host by Khlebtsov [17].
 Whereas the dipole MLWA continues to overlie with the exact Mie theory results for $a\lesssim 25$ nm irrespective of $\bar{\kappa}$, the results of the $\bar{\kappa}$-PT begin to deviate from those of the dipole MLWA for $\bar{\kappa}\gtrsim 0.05$.
-This is to be expected, because the $\bar{\kappa}$-PT is a first order perturbation theory of the dipole MLWA in $\bar{\kappa}$.
+An extreme case of such a deviation is provided for Ag sphere with $a=25$ nm and $\bar{\kappa}=\pm 0.1$ shown in Fig.
+4(f),(h).
+Whereas the dipole MLWA nearly overlies the exact result showing a positive (negative) extinction for $\bar{\kappa}=0.1$ ($\bar{\kappa}=-0.1$), the $\bar{\kappa}$-PT in the respective cases indicates a negative (positive) extinction.
+This is to be expected, because the $\bar{\kappa}$-PT is a first order perturbation theory of the dipole MLWA in $\bar{\kappa}$ and is expected to eventually break down above a certain threshold value of $|\bar{\kappa}|$.
+Negative extinction shown in Fig.
+4(h) is an indication of that particle losses have been more than compensated by the gain medium - a precursor of lasing action [37, 38, 39].
 Surprisingly, the smaller the particle, the greater the deviation of the $\bar{\kappa}$-PT relative to the dipole MLWA for $\bar{\kappa}\gtrsim 0.05$ (see Fig.
-S1 and S2 of ESI).
+S1 and S2 in Supplement 1).
 The origin of this behaviour is that $D(\varepsilon_{t})$ is typically small in a proximity of a LSPR (cf.
-the generalized Fröhlich condition eqn (5)).
-Whereas the first term in eqn (9) is of the order $1/D(\varepsilon_{t})$, the term proportional to $\bar{\kappa}$ is of the order $1/D^{2}(\varepsilon_{t})$.
+the generalized Fröhlich condition Eq.
+(5)).
+Whereas the first term in Eq.
+(9) is of the order $1/D(\varepsilon_{t})$, the term proportional to $\bar{\kappa}$ is of the order $1/D^{2}(\varepsilon_{t})$.
 However, when the size parameter $x$ increases, $|D(\varepsilon_{t})|$ at a LSPR decreases (due to a larger separation from the complex zero), and, as illustrated in Fig.
-5, $\bar{Q}_{\bar{\kappa}}$, defined as the ratio of the PT correction to the leading term on the rhs of eqn (9), gradually decreases, whereby the first-order $\bar{\kappa}$-PT begins to approximate the MLWA.
+5, $\bar{Q}_{\bar{\kappa}}$, defined as the ratio of the PT correction to the leading term on the rhs of Eq.
+(9), gradually decreases, whereby the first-order $\bar{\kappa}$-PT begins to approximate the MLWA.
 Note also how the ratio $\bar{Q}_{\bar{\kappa}}$ increases with $\bar{\kappa}$, which, as expected, explains why the $\bar{\kappa}$-PT begins to deviate from the MLWA with increasing $\bar{\kappa}$.
@@
 Again the smaller the particle, the greater the deviation.
-Nevertheless, the Mie theory approximation with nonabsorbing host becomes the best approximation with increasing particle radius (e.g.
-$a\gtrsim 40$ nm and $\bar{\kappa}\leq 0.01$).
+Nevertheless, in a relatively weakly absorbing host the Mie theory approximation with nonabsorbing host becomes the best approximation.
 For example, the Mie theory in nonabsorbing host begins to overlie with the Mie theory in an absorbing host for $a\gtrsim 50$ nm and $\bar{\kappa}=0.01$ (see Fig.
-S1 of ESI).
+S1 in Supplement 1).
 A threshold radius for which it happens increases with $\bar{\kappa}$.
@@
 This is demonstrated also in Fig.
-6 showing the effect of increasing the radius of spherical Ag nanoparticles in Poly(3-hexylthiophene) (P3HT) host medium, which was used in a number of recent studies 15, 37, 17, 19.
+6 showing the effect of increasing the radius of spherical Ag nanoparticles in Poly(3-hexylthiophene) (P3HT) host medium, which was used in a number of recent studies [15, 40, 17, 19].
 Unlike inherent dipole approximations, the Mie theory approximation with nonabsorbing host captures reasonably well both the dipole and quadrupole peaks.
@@
 4 (Fig.
-S5 of ESI) such as spasers 38, 39, 40.
+S5 in Supplement 1) such as spasers [37, 38, 39].
 Therefore, within the range of its validity, the first order $\bar{\kappa}$-PT allows one to both intuitively and analytically understand the mechanisms of host dissipation or gain on the extinction efficiency of a plasmonic nanosphere.
-Fig.
-7: The effect $\eta_{\rm eff}$ of the host absorption (left column) and the gain host medium (right column) on $Q_{\rm ext}$ in the Mie theory (solid red line) and the $\bar{\kappa}$-PT (dashed dark green line) for Al, Ag, Au and Mg materials with different radii $a$ embedded in the host medium with $n_{2}^{\prime}=1.33$ and $\bar{\kappa}=0.001$ (left column) and $\bar{\kappa}=-0.001$ (right column).
+Figure 7: The effect $\eta_{\rm eff}$ of the host absorption (left column) and the gain host medium (right column) on $Q_{\rm ext}$ in the Mie theory (solid red line) and the $\bar{\kappa}$-PT (dashed dark green line) for Al, Ag, Au and Mg materials with different radii $a$ embedded in the host medium with $n_{2}^{\prime}=1.33$ and $\bar{\kappa}=0.001$ (left column) and $\bar{\kappa}=-0.001$ (right column).
 4 Discussion
-Rather exhaustive MLWA analysis in the absorbing case has been recently provided by Khlebtsov 17 by employing slightly different MLWA.
+First-principles far-field computations based on the general Lorenz–Mie theory showed that increasing absorption in the host medium broadens and suppresses plasmon resonances in the apparent extinction [14].
+Such a broadening and suppression of plasmon resonances is also present in all approximation shown in Figs.
+3 and  4.
+The effect of absorption in the host thus goes in the same direction as increasing absorption within the particle.
+Noteworthy, increasing gain in the host medium narrows plasmon resonances and increases their amplitude in the apparent extinction as demonstrated in Fig.
+4 (cf.
+also Fig.
+S3).
+This goes along physical intuition.
+Contrary to the apparent extinction, in the case of the inherent cross-sections a surrounding lossy medium was shown to narrow the plasmonic resonances, as well as increase their amplitude dramatically [15, 17].
+The dipole MLWA can, in principle, be used up to the laser threshold associated with the so-called spectral singularity of non-Hermitian models.
+Indeed, as it has been verified both experimentally and theoretically more than 15 years ago, a description of gain media consisting in changing the sign of extinction is a valid approximation up to the laser threshold [41].
+This could be particularly important for applications in spasers and nano-lasers.
+For the sake of completeness, rather exhaustive MLWA analysis in the absorbing case has been recently provided by Khlebtsov [17] by employing slightly different MLWA.
 The study of ref.
@@
 The agreement for Au and Ag (Al and Mg) nanoparticles, slightly better in the case of Au than Ag, continues to be acceptable up to $a\sim 50$ nm ($a\sim 40$ nm), and can be used, at least qualitatively, up to $a\sim 70$ nm ($a\sim 50$ nm).
+Expanding the discussion to the case of larger nanoparticles with $a>70$ nm almost certainly requires consideration of MLWA for higher order multipoles [31, 34].
 We developed within the dipole MLWA a first order analytic perturbation theory (PT) around a nonabsorbing host in a normalized extinction coefficient $\bar{\kappa}$ and investigated its properties.
-It was shown that, in a suitable parameter range, the $\bar{\kappa}$-PT can reliably capture the effect of host absorption or gain on the extinction efficiency of a plasmonic nanosphere.
-Considering growing interest in light-matter interactions, we expect that our results will help in designing optimal systems comprising plasmonic nanoparticles embedded in suitable dissipative or gain media for various applications.
-Author Contributions
-Conflicts of interest
+It was shown that, in a suitable parameter range, the $\bar{\kappa}$-PT can reliably isolate and capture the effect of host absorption or host gain on the overall extinction efficiency of spherical plasmonic nanoparticles.
+Considering growing interest in light-matter interactions, we expect that our results will help in designing optimal systems comprising plasmonic nanoparticles embedded in suitable dissipative or gain media for various applications, such as photothermal therapy [42, 43], spasers and other active medium devices [37, 38, 39].
+6 Funding
+The analytical study of absorbing media is funded by Russian Science Foundation grant № 23-19-00511.
+7 Declaration of competing interest
 There are no conflicts to declare.
-Acknowledgements
-Electronic Supplementary Information
-The supplementary information includes Sec.
-S1 “Supplementary Figures”, where additional visual data are provided on apparent extinction efficiencies for further particle radii, host $n_{2}^{\prime}$, and normalized extinction coefficient $\bar{\kappa}$.
-In Sec.
-S2 “Complex Pole of the Dipole Contribution within the Dipole MLWA”, the predictive power of the complex pole knowledge on the LSPR position is investigated.
-Sec.
-S3 “Proof of that $\varepsilon_{E1}$ is the Complex Zero of the Denominator of $a_{1}$” presents mathematical proof of what is in its title.
-In Sec.
-S4 “The Imaginary Part of $\varepsilon_{E1}$ is Negative” proves the stated statement for any real $x$.
-In Sec.
-S5 “The Derivative $da_{1}/d\bar{\kappa}$” explicit calculations leading to eqn (7) are presented.
-Sec.
-S6 “A Previous Attempt to Formulate a Perturbation Theory (PT) in Terms of $\bar{\kappa}$” analyses an earlier failed theoretical attempt.
-Sec.
-S7 “Apparent Extinction Cross Section” reviews the absorption and scattering cross-sections within the MLWA.
-Finally, Sec.
-S8 “Conventional Scattering Theory” reviews some related aspects of conventional scattering theory.
-Notes and references
-  • Newton 1982 R. G. Newton, Scattering Theory of Waves and Particles, Springer Berlin Heidelberg, Berlin, Heidelberg, 1982.
-  • Bohren and Huffman 1998 C. F. Bohren and D. R. Huffman, Absorption and Scattering of Light by Small Particles, Wiley-VCH Verlag GmbH & Co. KGaA, 1998.
-  • Fuchs and Kliewer 1968 R. Fuchs and K. L. Kliewer, J. Opt. Soc. Am., 1968, 58, 319–330.
-  • Mundy et al. 1974 W. C. Mundy, J. A. Roux and A. M. Smith, J. Opt. Soc. Am., 1974, 64, 1593–1597.
-  • Chýlek and Pinnick 1979 P. Chýlek and R. G. Pinnick, Appl. Opt., 1979, 18, 1123–1124.
-  • Bohren and Gilra 1979 C. F. Bohren and D. P. Gilra, J. Colloid Interface Sci., 1979, 72, 215–221.
-  • Lebedev et al. 1999 A. N. Lebedev, M. Gartz, U. Kreibig and O. Stenzel, Eur. Phys. J. D., 1999, 6, 365–373.
-  • Sudiarta and Chylek 2001 I. W. Sudiarta and P. Chylek, J. Opt. Soc. Am. A, 2001, 18, 1275–1278.
-  • Yang et al. 2002 P. Yang, B.-C. Gao, W. J. Wiscombe, M. I. Mishchenko, S. E. Platnick, H.-L. Huang, B. A. Baum, Y. X. Hu, D. M. Winker, S.-C. Tsay and S. K. Park, Appl. Opt., 2002, 41, 2740–2759.
-  • Fu and Sun 2006 Q. Fu and W. Sun, J. Quant. Spectrosc. Radiat. Transf., 2006, 100, 137–142.
-  • Mishchenko 2007 M. I. Mishchenko, Opt. Express, 2007, 15, 13188–13201.
-  • Mishchenko et al. 2017 M. I. Mishchenko, G. Videen and P. Yang, Opt. Lett., 2017, 42, 4873–4876.
-  • Mishchenko and Yang 2018 M. I. Mishchenko and P. Yang, J. Quant. Spectrosc. Radiat. Transf., 2018, 205, 241–252.
-  • Mishchenko and Dlugach 2019 M. I. Mishchenko and J. M. Dlugach, Appl. Opt., 2019, 58, 4871–4877.
-  • Peck et al. 2019 R. L. Peck, A. G. Brolo and R. Gordon, J. Opt. Soc. Am. B, 2019, 36, F117–F122.
-  • Mishchenko et al. 2019 M. I. Mishchenko, M. A. Yurkin and B. Cairns, OSA Continuum, 2019, 2, 2362–2368.
-  • Khlebtsov 2021 N. G. Khlebtsov, Phys. Chem. Chem. Phys., 2021, 23, 23141–23157.
-  • Dong et al. 2021 J. Dong, W. Zhang and L. Liu, Opt. Express, 2021, 29, 7690–7705.
-  • Zhang et al. 2022 S. Zhang, J. Dong, W. Zhang, Minggang Luo and L. Liu, Opt. Lett., 2022, 47, 5577–5580.
-  • Mishchenko and Dlugach 2019 M. I. Mishchenko and J. M. Dlugach, OSA Continuum, 2019, 2, 3415–3421.
-  • Videen and Sun 2003 G. Videen and W. Sun, Appl. Opt., 2003, 42, 6724–6727.
-  • Meier and Wokaun 1983 M. Meier and A. Wokaun, Opt. Lett., 1983, 8, 581–583.
-  • Zeman and Schatz 1984 E. J. Zeman and G. C. Schatz, Dynamics on Surfaces, Dordrecht, 1984, pp. 413–424.
-  • Zeman and Schatz 1987 E. J. Zeman and G. C. Schatz, J. Phys. Chem., 1987, 91, 634–643.
-  • Kelly et al. 2003 K. L. Kelly, E. Coronado, L. L. Zhao and G. C. Schatz, J. Phys. Chem. B, 2003, 107, 668–677.
-  • Kuwata et al. 2003 H. Kuwata, H. Tamaru, K. Esumi and K. Miyano, Appl. Phys. Lett., 2003, 83, 4625–4627.
-  • Moroz 2009 A. Moroz, J. Opt. Soc. Am. B, 2009, 26, 517–527.
-  • Zorić et al. 2011 I. Zorić, M. Zäch, B. Kasemo and C. Langhammer, ACS Nano, 2011, 5, 2535–2546.
-  • Massa et al. 2013 E. Massa, S. A. Maier and V. Giannini, New J. Phys., 2013, 15, 063013.
-  • Le Ru et al. 2013 E. C. Le Ru, W. R. C. Somerville and B. Auguié, Phys. Rev. A, 2013, 87, 012504.
-  • Schebarchov et al. 2013 D. Schebarchov, B. Auguié and E. C. Le Ru, Phys. Chem. Chem. Phys., 2013, 15, 4233–4242.
-  • Januar et al. 2020 M. Januar, B. Liu, J.-C. Cheng, K. Hatanaka, H. Misawa, H.-H. Hsiao and K.-C. Liu, J. Phys. Chem. C, 2020, 124, 3250–3259.
-  • Rasskazov et al. 2020 I. L. Rasskazov, P. S. Carney and A. Moroz, Opt. Lett., 2020, 45, 4056–4059.
-  • Rasskazov et al. 2021 I. L. Rasskazov, V. I. Zakomirnyi, A. D. Utyushev, P. S. Carney and A. Moroz, J. Phys. Chem. C, 2021, 125, 1963–1971.
-  • McPeak et al. 2015 K. M. McPeak, S. V. Jayanti, S. J. P. Kress, S. Meyer, S. Iotti, A. Rossinelli and D. J. Norris, ACS Photonics, 2015, 2, 326–333.
-  • Palm et al. 2018 K. J. Palm, J. B. Murray, T. C. Narayan and J. N. Munday, ACS Photonics, 2018, 5, 4677–4686.
-  • Peck et al. 2021 R. Peck, A. Khademi, J. Ren, S. Hughes, A. G. Brolo and R. Gordon, Phys. Rev. Research, 2021, 3, 013014.
-  • Bergman and Stockman 2003 D. J. Bergman and M. I. Stockman, Phys. Rev. Lett., 2003, 90, 027402.
-  • Lawandy 2004 N. M. Lawandy, Appl. Phys. Lett., 2004, 85, 5040–5042.
-  • Veltri et al. 2016 A. Veltri, A. Chipouline and A. Aradian, Sci. Rep., 2016, 6, 33018.
+8 Data availability
+Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
+9 Supplemental document
+See Supplement materials for supporting content.